2. Find the remainder when r -ax2 + 6% - a UJ divided by
z - a.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
When p(x) = r -ax2 + & - a is divided by (x - a), the remainder is p(a). :. p(a) = a3 - a3 + 6a - a = 5c.